Rocky Guide Service provides guided 1-5 day hiking tours throughout the Rocky Mountains. Wilderness Tours hires Rocky to leadRequired: 1. to 3. Prepare the journal entries to record the transactions above. (If no entry is required for a transaction/eplease help!!!!
Answer #1

solution: No Date General journal 1 Jul-15 account receivable = 10*3000 bonus receivable =10*300*0.3 service revenue debit cr